OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation x squared over sixteen minus y squared over forty eight = 1

OpenStudy (tkhunny):

\(\dfrac{x^{2}}{16} - \dfrac{y^{2}}{48} = 1\) Cneter is (0,0), right? Vertices are in the x-direction a distance of 4. Why? Fake Vertices are in the y-direction a distance of \(4\sqrt{3}\). Why? What is the relationship between a, b, and c for an hyperbola?

OpenStudy (anonymous):

im sorry, this was the wrong question. could you help me with another thats similar to this?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation quantity x plus 5 squared divided by 36 minus the quantity of y plus 1 squared divided by 64 equals 1.

OpenStudy (tkhunny):

It's almost exactly the same written in this standard form. You tell me where the center is.

OpenStudy (anonymous):

(5,1) ?

OpenStudy (tkhunny):

No good. (-5,-1) Get used to it being (x-h) and (y-k)

OpenStudy (anonymous):

oh okay. so to find the foci i use (h+c,k) right? what would c be?

OpenStudy (tkhunny):

Don't get ahead of your self. Let's use the information we have. With +x^2 and -y^2, the vertices will be in the x0direction. Do you believe this?

OpenStudy (anonymous):

Yes, the hyperbola will be opening left/right

OpenStudy (tkhunny):

Okay, now let's look at a. a = 6. Do you see this?

OpenStudy (anonymous):

yes because the sq rt of 36 is 6

OpenStudy (tkhunny):

Perfect. This makes the vertices where?

OpenStudy (anonymous):

(1,-1)

OpenStudy (tkhunny):

That's one. Where is the other? You must go both ways, left and right.

OpenStudy (anonymous):

(-11,-1) ?

OpenStudy (tkhunny):

Awesome. There are two vertices. Okay, now find the value of b.

OpenStudy (anonymous):

sq rt of 64. so 8

OpenStudy (tkhunny):

Very good. This information will pin down the co-vertices. Where are they?

OpenStudy (anonymous):

are co-vertices the sam as foci?

OpenStudy (anonymous):

im not sure what co-vertices are

OpenStudy (tkhunny):

No. Co-vertices would be vertices if the hyperbola went the other way. I like to call them "fake vertices".

OpenStudy (tkhunny):

Center is (-5,-1) With b = 8, co-vertices are (-5,-9) and (-5,7) - just like the vertices, but in the other direction. Make sense?

OpenStudy (anonymous):

Oh okay, i get it

OpenStudy (tkhunny):

We're almost there. Given a and b, how do we find c?

OpenStudy (anonymous):

a^2 + b^2 = c^2 ?

OpenStudy (tkhunny):

Perfect. Calculate c.

OpenStudy (anonymous):

10

OpenStudy (tkhunny):

...and use that to find the foci!!

OpenStudy (anonymous):

so it would be (-15, -1), (5, -1)??

OpenStudy (tkhunny):

Yes! Excellent. There are two more things that are often asked. Asymptotes? Eccentricity? Do you need to supply these, or are we done?

OpenStudy (anonymous):

no, those arent in the question. thank you so much! i understand it way better now.

OpenStudy (tkhunny):

Perfect. When you get to those last two, come on back and we can drag through that, too.

OpenStudy (tkhunny):

Note: The real trick, here, was just to proceed one thing at a time. There is a lot of stuff. Just plod through it and don't try to jump to the end. :-)

OpenStudy (anonymous):

yes, it was much easier step by step